If country A has capital (K)=900, Labor (L)= 500, and country B has K=800 and L=400, which country should specialize and export capital intensive products?


1
Expert's answer
2021-04-22T12:23:32-0400

Country A autarky prices for:

Capital=K/L

K=900/500=1.8L

L=500/900=0.5k

Country B:

K=800/400=2L

L=400/800=0.25k

Thus, Country A should specialize and export capita intensive products because it uses less units of capital than co untry B